OpenStudy (anonymous):

Use an identity to solve the following equation on the interval [0,2pi) cos2x=cosx-1

OpenStudy (anonymous):

\[\cos 2x=\cos x-1,1+\cos 2 x=\cos x\] \[2 \cos ^2x=\cos x,\cos~ x ~\left( 2 \cos x-1 \right)=0\] solve

OpenStudy (anonymous):

Can you help elaborate more? Do you then use the unit circle or do you find the zeros and ones for cosx?

OpenStudy (anonymous):

\[\cos x=0=\cos \frac{ \pi }{ 2 } ,\cos \frac{ 3 \pi }{ 2 },x=?\]

OpenStudy (anonymous):

\[2 \cos x=1,\cos x=\frac{ 1 }{ 2 }=\cos \frac{ \pi }{ 3 },\cos \left( 2 \pi-\frac{ \pi }{ 3 } \right)\] x=?

OpenStudy (anonymous):

Okay, I got it! Thank you(:

OpenStudy (anonymous):

yw